In the class, we have learned a formula for the end-to-end delay of sending one packet of length L over 2 links of transmission rate R, (i.e., 2L/R). Give a formula for the total end-to-end delay of sending P packets of length L over N links of transmission rate R.
Answer #1

For one packet it is (NL)/R Ignoring queuing and processing delays

For P packets we will have

At t = NL/R, the first packet reaches the destination

At t = NL/R + L/R, the second packet reaches the destination because of the transmission delay of the first packet to the second link

Similarly…

At t = NL/R + (P-1)L/R , the last packet reaches the destination

Total end to end delay = (N+P-1)L/R

Do give a thumbs up and in case there are doubts leave a comment.